Two cities are 405 miles apart. A car leaves one of the cities traveling toward the second city at 60mph. At the same time, a bus leaves the second city bound for the first city at 75mph. How long will it take them to meet?
Answer by mananth(16949)

Two cities are 405 miles apart. A car leaves one of the cities traveling toward the second city at 60mph. At the same time, a bus leaves the second city bound for the first city at 75mph. How long will it take them to meet?
Let the distance traveled by the car when they meet be x. Speed of car is 60 mph
The bus would have traveled 405 -x miles . speed of bus is 75mph
Time taken by car when they meet will be x/60
Time taken by bus when they meet will be 405-x / 75
So x/60 = 405-x / 75
75x = 60 ( 405-x)
75x= 24300 -60x
75x+60x=24300
x= 24300 / 135
x= 180
180 /60 =3 hours
They meet after 3 hours.